OpenStudy (anonymous):

solve and check fan and medal C/3 =6 1/7 A.2 1/21 B.2 1/7 C.18 1/7 D.18 3/7

OpenStudy (anonymous):

c/3= 6 1/7 (change to improper) (7*6 +1)

OpenStudy (anonymous):

c/3 = 43/7

OpenStudy (anonymous):

c/3 *3 = 43/7 * 3 (*3 to both side to isolate c)

OpenStudy (anonymous):

c= 47/3 * 3

OpenStudy (anonymous):

c= 129/7 (change to mix)

OpenStudy (anonymous):

@osamabinswaggin tanks u soooo much

OpenStudy (anonymous):

c= 18 3/7

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

medal?

OpenStudy (anonymous):

there i gave u a medal and fan